# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v2

@@ IV Methods
To study the performance of our proposed X3Z3 code, we perform quantum memory experiment simulations, with varying degrees of noise bias $\eta$, of this code along with three other codes: CSS, P6 and XYZ2 Floquet codes. The noise bias $\eta=p_{Z}/(p_{X}+p_{Y})$ is defined as the ratio of the $Z$-error $p_{Z}$ to other errors where the total physical error rate is $p=p_{X}+p_{Y}+p_{Z}$. As the noise asymmetry increases, the noise bias $\eta$ increases from $\eta=0.5$, which corresponds to fully depolarising noise ($p_{X}=p_{Y}=p_{Z}=p/3$), to $\eta=\infty$, corresponding to a pure $Z$-biased noise ($p_{Z}=p$, and $p_{X}=p_{Y}=0$).
We use two different biased noise models which are the generalisations of the standard code-capacity and standard depolarising entangling measurement (SDEM3) [9, 15] noise models to biased noise channels. The code-capacity noise model is one in which only single-qubit errors are applied to all data qubits at every measurement subround. We note that, since for Floquet codes one QEC round consists of several measurement subrounds, the single-qubit noise channel in the code-capacity model is applied several times in one QEC round instead of just once as is the case with static codes. The code-capacity noise model is often used as a preferred initial noise model to study before going to a more involved model as its simplicity often offers insight into understanding the code performance.
On the other hand, the SDEM3 model is a more elaborate error model involving single-qubit noise channels after every single-qubit gate, measurement and reset, a two-qubit noise channel after every multi-Pauli product (MPP) parity measurement gate, and a classical flip after each measurement. As in Ref. [9], we assume that each of these error channels occurs with a total probability $p$. The SDEM3 noise model is therefore close to standard circuit-level noise and would be a more accurate description of a realistic noise channel, particularly in hardwares with native two-qubit measurements, for example, Majorana [56, 10, 57] and photonic [58, 17] architectures. To take into account biased noise for the MPP gates, we here generalise the SDEM3 noise model in Refs. [9, 15]. The details of the biased code-capacity and SDEM3 noise models can be found in Appendix B.
We simulate the circuits and generate the error syndromes using Stim [64] (example Stim circuits and threshold data are provided at Ref. [65]). To decode the error syndromes, we apply a minimum-weight perfect matching (MWPM) decoder, which is implemented using PyMatching [55]. Our simulations are run for codes with effective distances $d_{\mathrm{eff}}\equiv d/2=4,6,8,10,12$, where the effective distance $d_{\mathrm{eff}}$ is defined as the minimum number of faults under SDEM3 depolarising noise that produce a logical error. This is half of the distance $d$ of the code-capacity noise model. The calculated effective distance differs between noise models since in the SDEM3 noise model, two-qubit errors occurring after an MPP parity gate count as a single fault, while under the code-capacity noise model they would count as two faults. We choose lattices with periodic boundary conditions and sizes $L\times 3L/2$. Here, $L=d=2d_{\mathrm{eff}}$ is the lattice length which varies with respect to the code distance. We simulate the memory experiments for $3d_{\mathrm{eff}}$ QEC rounds, with each round containing 6 subrounds of check measurements.
V Results and Discussion
@@
Figure 7: Sub-threshold logical failure probability $p_{L}$ of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a) code-capacity and (b) SDEM3. Inset: Zoom-in subthreshold performance plots of the P6 (red) and XYZ2 (orange) codes. Results are calculated using (a) lattice size $20\times 30$ ($d=20$) for 30 QEC rounds with $p=0.72\%$ and (b) lattice size $24\times 36$ ($d_{\mathrm{eff}}=12$) for 36 QEC rounds with $p=0.5\%$. Each data point is averaged over $10^{6}-10^{8}$ Monte Carlo shots. For better visualisation, we fit all curves with quadratic splines.
Figure 6 shows the thresholds of all codes for various levels of noise bias $\eta$ calculated for (a) code-capacity and (b) SDEM3 noise models. Each threshold is obtained from the intersection of the logical failure probabilities $p_{L}$ vs physical error rate $p$ curves of different code distances $d_{\mathrm{eff}}=d/2=6,8,10$ and 12. The data of $p_{L}$ as a function of $p$ for each code and various noise-bias strengths, used to obtain the thresholds, are presented in Figs. 12-15 of Appendix C.
@@
It is interesting to note that for the fully depolarising code-capacity noise, all codes have the same threshold. This is because at every measurement subround under this noise model, all the above codes have two types of errors at each fault location (occurring with a total probability $32p/$ 3) that give rise to edges in the detector graph and one type (occurring with probability $p/3$) that produces hyperedges. For the CSS and X3Z3 codes, these hyperedges result from $Y$ Pauli errors. On the other hand, the hyperedge errors in the P6 and XYZ2 codes are those that anti-commute with the check measurements that occur just before and immediately after the errors [12, 22]. Therefore, the Pauli type of the hyperedge errors for the honeycomb codes varies between measurement subrounds. For example, $Z$ errors create hyperedges only when they occur between the $XX$ and $YY$ checks. In summary, all the above codes have the same performance for fully depolarising code-capacity noise because their (weighted) detector hypergraphs are all equivalent.
@@
The performance difference between the XYZ2 and P6 honeycomb codes can be understood from the distribution of hyperedge-like syndromes (namely, those with four triggered detectors) in the detector hypergraphs under the biased SDEM3 noise model. These hyperedges generally degrade the code’s performance when using a matching decoder since they must be decomposed into edges (see Ref. [12] for a detailed description of the decoding in a hyperbolic version of the honeycomb code). An infinitely biased SDEM3 noise model that includes Pauli noise but no measurement errors can already form hyperedges in the honeycomb codes’ detector hypergraphs. In particular, the P6 code can have hyperedges resulting from $Z$ errors occurring only after the red check measurement subrounds, while the hyperedges in XYZ2 code can result from certain $Z$ errors happening in any subround. While there is no difference in the overall number of hyperedges resulting from single-qubit $Z$ errors in the two hypergraphs, the different arrangement of the hyperedges due to the single-qubit errors in these two codes gives rise to different kinds of syndromes for the two-qubit $ZZ$ errors. Specifically, $ZZ$ errors can lead to hyperedge syndromes in the XYZ2 Floquet code but only edge-like syndromes for the P6 code (see Appendix D for details). As a result, in an infinitely biased regime, where Pauli errors after MPP gates are evenly distributed between $Z_{1}$, $Z_{2}$ and $Z_{1}Z_{2}$, there are more hyperedge syndromes triggered in the XYZ2 code than in the P6 code. This is the reason underlying the difference in the thresholds, as shown in Fig. 6(b), of the two honeycomb codes under the SDEM3 noise. Since honeycomb codes are prone to hyperedge errors, better performance of the honeycomb codes might be expected using a correlated MWPM decoder such as the one used in Ref. [8].
As a result of the shift of the thresholds with the noise bias, the subthreshold performance of the codes changes accordingly. As shown in Figs. 7(a) and 7(b), the subthreshold logical failure probability improves significantly for the X3Z3 code, while it becomes only slightly better for the honeycomb codes and deteriorates for the CSS code.
@@
Figure 8: Exponential suppression of the sub-threshold logical failure probabilities $p_{L}$ with respect to $d$ or $d_{\mathrm{eff}}$ for different Floquet codes: CSS (green), XYZ2 (orange), P6 (red), and X3Z3 (blue). Results are calculated for different noise models: (a,b) code-capacity and (c,d) SDEM3 noise models, with a physical error rate $p=0.55\%$. They are also computed using different bias strengths: (a,c) $\eta=1$ and (b,d) $\eta=99$. All curves can be fitted to an exponential decay function $f\propto\mathrm{exp}(-\gamma d)$ or $f\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$ where $\gamma$ depends on the bias strength $\eta$ and is an increasing function of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{5}-10^{9}$ shots.
@@
Besides thresholds, another quantity that is of important is the scaling of the code sub-threshold performance with the code distance. To show this, we plot the sub-threshold logical failure probability as a function of $d$ or $d_{\mathrm{eff}}$ for varying levels of noise bias in Fig. 8. As shown, the logical failure probabilities for all codes decrease exponentially with the code distance, i.e., $p_{L}\propto\mathrm{exp}(-\gamma d)$ or $p_{L}\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$. Among all codes presented, the X3Z3 code has the largest logical error suppression rate $\gamma$. At higher noise bias, this error suppression rate becomes significantly larger for the X3Z3 code, moderately increases for the two honeycomb codes and decreases for the CSS Floquet code. The reason is that as the noise bias increases, a fixed subthreshold physical error rate moves relatively with respect to the shifting threshold, so that it becomes much further below the threshold for the X3Z3 Floquet code, moves moderately away from the threshold for the two honeycomb codes and becomes closer to the threshold of the CSS code [see Figs. 6(a) and 6(b)].
@@
We also numerically simulate the performance of the X3Z3 Floquet code with a twisted periodic boundary condition. This twisted code has a pure $Z$-type logical operator with a length that scales quadratically with the code distance of the untwisted code. As a result, we find better thresholds and more favorable subthreshold performance for the twisted code over the untwisted one. While these performance benefits hold for the infinite-bias code-capacity noise, they do not apply to other noise biases and models tested (See Appendix E for details).
@@
The error suppression rate $\gamma$ is related to the error suppression factor $\Lambda$ introduced in Ref. [66] which is defined as the reduction factor in the logical failure probability as the code distance increases by 2. Mathematically, it is given by $\Lambda\equiv p_{L}(d)/p_{L}(d+2)=\mathrm{exp}(2\gamma)$ for the code-capacity noise and similarly for the SDEM3 noise model but with $d$ replaced by $d_{\mathrm{eff}}$. In Table 1 of Appendix C.2 we list the values of $\Lambda$ corresponding to the $\gamma$ values shown in Fig. 8.
VI No-go theorems for 1D decoding graphs of dynamical codes
Figure 9: An illustration of (a) Theorem 1 and (b) Theorem 2. We display examples of the $Z$-detector graphs (edges correspond to $Z$ errors and vertices to detectors) that are forbidden by these theorems, for (a) HDCs and (b) GDCs.
@@ VI No-go theorems for 1D decoding graphs of dynamical codes
We will start by formalising the above notions via:
Definition 1.
Definition 2.
Finally, we define dynamical codes in the following way:
Definition 3.
A honeycomb dynamical code (HDC), with “duration” $T$ and constant “initial(final)-time boundary offsets” $k_{i}$ ($k_{f}$), is a finite-depth measurement circuit acting on qubits of the honeycomb lattice (without boundary). The sets of measurements in the circuit, $\mathcal{M}_{t}$ ($t\in\{1,\ldots,T\}$), are composed of two-body Pauli measurements along coloured edges (either $r$, $g$, or $b$, for each time step) of the lattice, performed sequentially on a state stabilised by some group $\mathcal{S}$. The HDC obeys the following properties:
@@
    Overlapping measurements in consecutive time steps anti-commute and are supported on different edges,
@@
    Detectors (for time steps $t>k_{i}$) are associated with plaquettes in the lattice (they have support only on qubits around their associated plaquette),
@@
    If a single-qubit or two-qubit error occurring in time steps $k_{i}<t<T-k_{f}$ anti-commutes with future measurements, it is detectable, unless the two-qubit error is the same as an edge operator just measured or to be measured in the next time step.
  4. (d)
    All single-qubit errors in time steps $k_{i}<t<T-k_{f}$ have syndromes of weight $>1$.
Let us first comment on our definition of an HDC. The chosen properties are very natural. The anti-commutation of consecutive measurement subrounds ensures “local reversibility” of the code, so as to preserve the quantum information and its locality  [67] (although note that we do not require the code to have logical qubits in our definition). The requirement that consecutive sets of measurements are supported on different edges is not very restrictive: if two (anti-commuting) measurements act on the same edge in consecutive time steps, we can replace the second with a Clifford gate and commute that to the end of the circuit. This merely changes the bases of subsequent measurements without changing their (anti-)commutation. Since the Clifford gates are unimportant, we can ignore them and hence we end up with a shorter duration $T$. We restrict our attention to errors occurring in the temporal “bulk” of the code, $k_{i}<t<T-k_{f}$. This avoids complications due to errors occurring close to the initial and final-time boundaries where detectors here may be formed in different ways than the bulk detectors. Specifically, the detectors associated with errors at the initial and final time boundaries are obtained by comparing the edge or plaquette measurement values with the initial state and final read-out of the physical qubits, respectively. The exclusion of these final time boundary detectors is also required here since we do not include final read-out measurements in our definition of an HDC. In summary, we consider errors occurring in all “detection cells” [13] that are completed before time step $T$ and begin at time step $t\geq 1$, where a detection cell consists of all the spacetime points at which a detector can identify errors.
We will show that an HDC cannot even have a 1D-decodable ZDG, which implies that its decoding graph cannot be equivalent to a set of the much simpler repetition code’s graph. Note that a repetition code graph is not only 1D-decodable but also has a maximum degree of 2, which means there are no double edges between neighbouring vertices. Here, however, we allow for these double edges in our definition of 1D-decodability, and show that this more general property is also impossible. Such double edges naturally arise in Floquet codes, where they correspond to two-qubit undetectable errors (i.e., edge operators just measured or about to be measured).
Using the properties of the HDC as defined above, we now present the first main theorem of our paper (see Appendix F for a proof):
@@
A honeycomb dynamical code with duration more than $33k_{f}+k_{i}+$ 3 cannot have a 1D-decodable $Z$-detector graph, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets.
The above theorem implies that no matter how an HDC is bias-tailored, it likely can never achieve as high a threshold as the repetition code or the bias-tailored stabiliser codes. Indeed, for large $T$ ($T\gg k_{i},k_{f}$), there will be at least approximately $6$ times as many edges as vertices in the ZDG, and there is a subgraph with $3/2$ as many edges as vertices without any pair of vertices sharing more than one edge (see Appendix F). This means the graph must have many cycles with length at least three edges.
Definition 4.
A general dynamical code (GDC) is a finite-depth measurement circuit acting on qubits associated with vertices of a graph $G$. The sets of measurements $\mathcal{M}_{t}$ are two-body Pauli measurements associated with edges of $G$ (each qubit is in the support of one measurement in each time step), performed sequentially on a state stabilised by some group $\mathcal{S}$. A GDC obeys properties (a) and (d) of an HDC.
Theorem 2.
A GDC with a duration at least $22k_{f}+k_{i}+$ 2 (i.e., a constant), where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets, respectively, cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.
The proof of the theorem is given in Appendix F. Owing to this theorem, we do not expect to see thresholds in dynamical codes as high as 50%. The reason that static codes perform better than dynamical codes under biased noise is that they obey two separate symmetries under an infinitely-biased code-capacity noise model. There is a spatial symmetry that forces syndromes to appear in pairs along 1D strips of the lattice, but there is also a temporal symmetry: syndromes must appear in pairs within a given time slice. These two symmetries simplify the detector graph into a collection of disjoint repetition code decoding graphs, leading to a high performance (up to 50% threshold) of the static codes under biased noise. On the other hand, even for the case where measurement errors are not considered, GDCs necessarily break the temporal symmetry (for HDCs, this is a consequence of Lemma 1 in Appendix F), leading to edges between vertices in different time slices of the detector graph (see Fig. 3). We thus expect that such codes can only possess a single symmetry under an infinitely-biased code-capacity noise model, as the X3Z3 Floquet code does.
VII Conclusions and outlook

### v3

@@ II.7 Memory experiment simulation
To study the performance of our proposed X3Z3 code, we perform quantum memory experiment simulations under two biased noise models. These noise models are the generalisations of the standard code-capacity and standard depolarising entangling measurement (SDEM3) noise models (see the Methods section for details of the noise models). We simulate the X3Z3 code along with three other codes: CSS, P6 and XYZ2 Floquet codes, with varying degrees of noise bias $\eta$ (see the Methods section for details of our numerical simulations). The noise bias $\eta=p_{Z}/(p_{X}+p_{Y})$ is defined as the ratio of the $Z$-error $p_{Z}$ to other errors where the total physical error rate is $p=p_{X}+p_{Y}+p_{Z}$. As the noise asymmetry increases, the noise bias $\eta$ increases from $\eta=0.5$, which corresponds to fully depolarising noise ($p_{X}=p_{Y}=p_{Z}=p/3$), to $\eta=\infty$, corresponding to a pure $Z$-biased noise ($p_{Z}=p$, and $p_{X}=p_{Y}=0$).
@@
Figure 7: Sub-threshold logical failure probability $p_{L}$ of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a) code-capacity and (b) SDEM3. Inset: Zoom-in subthreshold performance plots of the P6 (red) and XYZ2 (orange) codes. Results are calculated using (a) lattice size $20\times 30$ ($d=20$) for 30 QEC rounds with $p=0.72\%$ and (b) lattice size $24\times 36$ ($d_{\mathrm{eff}}=12$) for 36 QEC rounds with $p=0.5\%$. Each data point is averaged over $10^{6}-10^{8}$ Monte Carlo shots. For better visualisation, we fit all curves with quadratic splines. For each of the noise models, we choose the largest code distance and the smallest subthreshold physical error rate that can be simulated given our computational resources.
Thresholds
Figure 6 shows the thresholds of all codes for various levels of noise bias $\eta$ calculated for (a) code-capacity and (b) SDEM3 noise models. Each threshold is obtained from the intersection of the logical failure probabilities $p_{L}$ vs physical error rate $p$ curves of different code distances $d_{\mathrm{eff}}\equiv d/2=6,8,10$ and 12. Here, the effective distance $d_{\mathrm{eff}}$ is defined as the minimum number of faults under SDEM3 depolarising noise that produce a logical error. This is half of the distance $d$ of the code-capacity noise model. For the methods and plots used to determine the thresholds, see the Methods section and Supplementary Figs. 5-8 of the Supplementary Information [61]. We also provide the threshold data in Ref. [69].
@@
It is interesting to note that for the fully depolarising code-capacity noise, all codes have the same threshold. This is because at every measurement subround under this noise model, all the above codes have two types of errors at each fault location (occurring with a total probability $32p/$ 3) that give rise to edges in the detector graph and one type (occurring with probability $p/3$) that produces hyperedges. For the CSS and X3Z3 codes, these hyperedges result from $Y$ Pauli errors. On the other hand, the hyperedge errors in the P6 and XYZ2 codes are those that anti-commute with the check measurements that occur just before and immediately after the errors [12, 22]. Therefore, the Pauli type of the hyperedge errors for the honeycomb codes varies between measurement subrounds. For example, $Z$ errors create hyperedges only when they occur between the $XX$ and $YY$ checks of the honeycomb codes. In summary, all the above codes have the same performance for fully depolarising code-capacity noise because their (weighted) detector hypergraphs are all equivalent.
@@
Figure 8: Exponential suppression of the sub-threshold logical failure probabilities $p_{L}$ with respect to code distance. We use code distances $d$ or $d_{\mathrm{eff}}$ depending on the noise model and compare different Floquet codes: CSS (green), XYZ2 (orange), P6 (red), and X3Z3 (blue). Results are calculated for different noise models: (a,b) code-capacity and (c,d) SDEM3 noise models, with a subthreshold physical error rate $p=0.55\%$, which is small enough, yet can still be simulated using our computational resources up to code distance $d=24$ or $d_{\mathrm{eff}}=12$, for all the different codes and parameters corresponding to different curves in the plots. The same physical error rates are chosen for all curves such that the subthreshold performance under different noise bias strengths and noise models can be compared. Plots are computed using two different bias strengths; one representing noise near the depolarizing regime: (a,c) $\eta=1$ and the other representing noise in the strongly dephasing regime: (b,d) $\eta=99$. All curves can be fitted to an exponential decay function $f\propto\mathrm{exp}(-\gamma d)$ or $f\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$ where $\gamma$ depends on the bias strength $\eta$ and is an increasing function of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{5}-10^{9}$ shots.
The performance difference between the XYZ2 and P6 honeycomb codes can be understood from the distribution of hyperedge-like syndromes (namely, those with four triggered detectors) in the detector hypergraphs under the biased SDEM3 noise model. These hyperedges generally degrade the code’s performance when using a matching decoder since they must be decomposed into edges (see Ref. [12] for a detailed description of the decoding in a hyperbolic version of the honeycomb code). An infinitely phase-biased SDEM3 noise model that includes Pauli noise but no measurement errors can already form hyperedges in the honeycomb codes’ detector hypergraphs. In particular, the P6 code can have hyperedges resulting from $Z$ errors occurring only after the red check measurement subrounds, while the hyperedges in XYZ2 code can result from certain $Z$ errors happening in any subround. While there is no difference in the overall number of hyperedges resulting from single-qubit $Z$ errors in the two hypergraphs, the different arrangement of the hyperedges due to the single-qubit errors in these two codes gives rise to different kinds of syndromes for the two-qubit $ZZ$ errors. Specifically, $ZZ$ errors can lead to hyperedge syndromes in the XYZ2 Floquet code but only edge-like syndromes for the P6 code (see Supplementary Sec. IV of the Supplementary Information [61] for details). As a result, in an infinitely phase-biased regime, where Pauli errors after MPP gates are evenly distributed between $Z_{1}$, $Z_{2}$ and $Z_{1}Z_{2}$, there are more hyperedge syndromes triggered in the XYZ2 code than in the P6 code. This is the reason underlying the difference in the thresholds, as shown in Fig. 6(b), of the two honeycomb codes under the SDEM3 noise. Since honeycomb codes are prone to hyperedge errors, better performance of the honeycomb codes might be expected using a correlated MWPM decoder such as the one used in Ref. [8].
II.8 Subthreshold performance
As a result of the shift of the thresholds with the noise bias, the subthreshold performance of the codes changes accordingly. As shown in Figs. 7(a) and 7(b), the subthreshold logical failure probability improves significantly for the X3Z3 code, while it becomes only slightly better for the honeycomb codes and deteriorates for the CSS code. In general, one expects the subthreshold performance of the codes to improve by increasing the code distance along the direction where the biased errors most likely form logical strings. For example, for the $Z$ biased noise considered here, it would be by increasing the vertical code distance $d_{V}$. To this end, we study the subthreshold performance of elongated CSS and X3Z3 Floquet codes with $d_{V}>d_{H}$, where $d_{V}$ and $d_{H}$ are the vertical and horizontal code distances, respectively. The details are given in Supplementary Sec. V of the Supplementary Information [61].
@@ II.8 Subthreshold performance
Besides thresholds, another quantity of importance is the scaling of the code sub-threshold performance with the code distance. We plot the sub-threshold logical failure probability as a function of $d$ or $d_{\mathrm{eff}}$ for varying levels of noise bias in Fig. 8. As shown, the logical failure probabilities for all codes decrease exponentially with the code distance, i.e., $p_{L}\propto\mathrm{exp}(-\gamma d)$ or $p_{L}\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$. Among all codes presented, the X3Z3 code has the largest logical error suppression rate $\gamma$. At higher noise bias, this error suppression rate becomes significantly larger for the X3Z3 code, moderately increases for the two honeycomb codes and decreases for the CSS Floquet code. The reason is that as the noise bias increases, a fixed subthreshold physical error rate moves relatively with respect to the shifting threshold, so that it becomes much further below the threshold for the X3Z3 Floquet code, moves moderately away from the threshold for the two honeycomb codes and becomes closer to the threshold for the CSS code [see Figs. 6(a) and 6(b)].
@@
We also numerically simulate the performance of the X3Z3 Floquet code with a twisted periodic boundary condition. This twisted code has a pure $Z$-type logical operator with a length that scales quadratically with the code distance of the untwisted code. As a result, we find better thresholds and more favorable subthreshold performance for the twisted code over the untwisted one. While these performance benefits hold for the infinite-bias code-capacity noise, they do not apply to other noise biases and models tested (See Supplementary Sec. VI of the Supplementary Information [61] for details).
While in this paper, we focus on Floquet codes with periodic boundary conditions, we expect similar qualitative performance of the Floquet codes under open boundary conditions. As shown in Ref. [9] for the honeycomb code with depolarizing noise, the boundary conditions (whether periodic or open) do not significantly impact the threshold and subthreshold performance of the code. This should be expected even more so for the CSS and X3Z3 Floquet codes, because the boundary condition does not affect the measurement schedule for the codes. We leave the simulations of the codes under open boundary conditions to future works.
@@
The error suppression rate $\gamma$ is related to the error suppression factor $\Lambda$ introduced in Ref. [70] which is defined as the reduction factor in the logical failure probability as the code distance increases by 2. Mathematically, it is given by
$\Lambda\equiv\frac{p_{L}(d)}{p_{L}(d+2)}=e^{2\gamma}$ (4)
for the code-capacity noise and similarly for the SDEM3 noise model, but with $d$ replaced by $d_{\mathrm{eff}}$. For a noise model where the error is characterised by a single error rate $p$ and threshold value $p_{\mathrm{th}}$, when $p\ll p_{\mathrm{th}}$, QEC suppresses the logical error exponentially as the code distance increases, i.e., $p_{L}\approx(p/p_{\mathrm{th}})^{d/2}$. Since $\Lambda\equiv p_{L}(d)/p_{L}(d+2)$, we then have $\Lambda\approx p_{\mathrm{th}}/p$. At the thresholds, we have $\Lambda(p_{\mathrm{th}})\approx 1$. The value of $\Lambda$ increases as the physical error rate $p$ moves further below the threshold $p_{\mathrm{th}}$. For physical error rate below the threshold, $\Lambda>1$ and larger $\Lambda$ means greater error correction. Since the threshold depends on the noise bias $\eta$, $\Lambda$ depends on both $\eta$ and $p$. In Table 1 we list the values of $\Lambda$ corresponding to the $\gamma$ values shown in Fig. 8.
Code-capacity SDEM3
$\eta=1$ $\eta=99$ $\eta=1$ $\eta=99$
CSS 1.94 1.39 1.74 1.37
XYZ2 2.27 2.28 1.11 1.15
P6 2.31 2.37 1.12 1.24
X3Z3 2.49 4.26 1.94 3.42
Table 1: Error suppression factors $\Lambda$ for different Floquet codes. These are calculated using two different noise models, code-capacity and SDEM3, with varying levels of noise bias $\eta$. Results are calculated with a physical error rate $p=0.55\%$. The lambda factor is related to the error suppresion rate $\gamma$ shown in Fig. 8.
II.9 Performance optimality of X3Z3 Floquet code and no-go theorems
@@
We argue that due to the above limiting constraint, the symmetry of the X3Z3 Floquet code has likely rendered its performance close to optimal under a matching decoder, despite the fact that even for the most favorable case of infinitely phase-biased code-capacity noise, its decoding can only be reduced to at most a series of disjoint 2D (planar) graphs. We formalise this by proving two no-go theorems for the decoding graphs of dynamical codes. We begin with the following informal statement of our first theorem:
Theorem 1.
(Informal) A dynamical code (not necessarily a Floquet code) on the honeycomb lattice operated over a sufficiently large number of time steps cannot have a 1D decoding graph under an infinitely biased code-capacity noise model, so long as it obeys the following properties:
@@
    At each time step, two-body Pauli operators on edges of a given colour are measured. Measurements in consecutive time steps occur on edges of different colours and anti-commute.
@@
    Detectors are supported on plaquettes.
@@
    All non-trivial errors are detectable and produce syndromes of weight $>1$.
In principle, a 1D graph [such as the one shown in Fig. 9(a)] could still result in better code performance than a general planar decoding graph. The fact that this type of graph is not possible suggests that planar graph decoding is already optimal for dynamical codes on the honeycomb lattice.
Figure 9: Illustrations of the no-go theorems. We display examples of the decoding graphs (under an infinitely biased code-capacity noise model; edges correspond to $Z$ errors and vertices to detectors) that are forbidden by (a) Theorem 1 for dynamical codes on the honeycomb lattice and (b) Theorem 2 for general dynamical codes.
To generalise this result beyond codes defined on the honeycomb lattice, we provide the following theorem for general dynamical codes built from two-qubit parity measurements.
@@
(Informal) A dynamical code (not necessarily a Floquet code) operated over a sufficiently long time scale cannot have a decoding graph, under an infinitely biased code-capacity noise model, equivalent to that of a collection of repetition codes [see Fig. 9(b)], so long as: all measurements are two-qubit parity measurements, each qubit is in the support of one measurement in each time step, overlapping measurements in consecutive time steps anti-commute, and errors produce syndromes of weight $>1$.
This theorem shows that (given reasonable and typical assumptions) the decoding graphs of general dynamical codes cannot be decomposed into a collection of 1D repetition-code decoding graphs, which could otherwise allow them to achieve higher thresholds.
We provide the formal statements of the above two theorems and their proofs in the Methods section. We display the decoding graphs prohibited by the above two theorems in Fig. 9.
III Discussion